# Revisiting LLM Evaluation through Mechanism Interpretability: a New Metric and Model Utility Law

### v1

@@ B.1 Utility Law: the Relationship Between MUI and Performance
Llama-2-7B-Chat 25.8 / 3.8 5.4 / 4.9 60.0 / 5.4 16.4 / 0.9 24.8 / 1.4 26.2 / 8.5 48.2 / 15.8
CodeLlama-7B-Instruct 23.0 / 4.5 6.0 / 7.0 49.5 / 8.7 34.1 / 1.8 44.7 / 2.9 21.8 / 8.3 41.4 / 24.5
Qwen2.5-7B-Instruct 84.5 / 1.9 61.9 / 2.8 86.8 / 4.2 71.3 / 0.7 62.1 / 1.0 66.0 / 5.3 71.3 / 14.5
Qwen2.5-Coder-7B-Instruct 79.6 / 3.7 49.5 / 6.2 83.8 / 9.4 75.0 / 1.5 65.9 / 2.4 57.8 / 11.0 65.6 / 30.3
Qwen2.5-Math-7B-Instruct 92.1 / 1.9 76.7 / 3.3 66.3 / 3.8 53.0 / 0.7 47.9 / 1.1 29.1 / 2.9 51.2 / 8.5
@@
Llama-2-7B-Chat 25.8 / 3.8 5.4 / 4.9 60.0 / 5.4 16.4 / 0.9 24.8 / 1.4 26.2 / 8.5 48.2 / 15.8
Llama-Code-Leakage 22.2 / 3.7 5.2 / 4.5 60.7 / 4.8 35.4 / 2.2 49.7 / 1.2 23.6 / 7.7 48.1 / 14.0
Llama-Math-Leakage 39.0 / 4.2 8.2 / 6.2 49.5 / 7.6 5.5 / 1.0 8.6 / 1.6 25.2 / 6.4 43.9 / 23.8
Llama-3.1-8B-Instruct 86.5 / 2.7 48.1 / 3.8 81.9 / 3.8 62.2 / 1.2 57.9 / 2.2 65.4 / 7.7 64.4 / 14.5
Llama3.1-Code-Leakage 86.6 / 3.1 40.2/ 4.5 81.2 / 5.0 81.7 / 1.2 69.1 / 1.9 48.1 / 6.9 66.3 / 15.1
Llama3.1-Math-Leakage 89.8 / 3.0 47.4 / 4.1 81.6 / 4.7 54.9 / 1.3 50.5 / 2.2 48.9 / 6.3 66.1 / 14.4
Qwen2.5-7B-Instruct 84.5 / 1.9 61.9 / 2.8 86.8 / 4.2 71.3 / 0.7 62.1 / 1.0 66.0 / 5.3 71.3 / 14.5
Qwen2.5-Code-Leakage 78.6 / 2.0 42.5 / 2.8 88.9 / 3.5 85.4 / 1.9 73.5 / 1.1 56.8 / 5.3 56.8 / 11.7
Qwen2.5-Math-Leakage 98.1 / 2.4 84.6 / 3.2 57.4 / 3.5 2.2 / 1.3 1.8 / 0.8 31.4 / 4.9 44.7 / 11.1
@@ B.6.3 Impact of Different Neuron Interpretation Methods
To demonstrate that our analysis is both generalizable and inclusive, we expand to incorporate various interpretation methods. Continuing from our previous threshold selections B.6.3, we conduct analyses using these methods to evaluate changes in the Model Utilization Index (MUI), with a particular focus on the consistency of MUI trends for different model. For some threshold values that do not integrate effectively, we randomly selected a threshold value within the testing range to facilitate a quantitative comparison. The results, which encompass the full parameter range, are detailed in the Figure 27. After applying different interpretation methods, we found that the MUI derived from various approaches (Table 13 and Table 14) correlates well with our experiment selections (Table 11), achieving an average Spearman correlation coefficient of 0.723. This indicates that the overall trend is relatively consistent that model capability and MUI is in a reverse Relationship. However, due to differences in the definition of importance, some models, such as Qwen2.5, exhibit a higher MUI when using activation-based methods compared to Llama3.1.
@@
Vicuna-7B 11.9 / 5.0 4.0 / 5.1 54.0 / 7.1 13.4 / 1.1 23.6 / 1.7 28.6 / 7.9 45.3 / 24.4
Llama-2-7B-Chat 25.8 / 3.6 5.4 / 4.6 60.0 / 5.0 16.4 / 0.9 24.8 / 1.3 26.2 / 7.9 48.2 / 14.9
Qwen1.5-7B-Chat 58.7 / 2.8 17.9 / 3.4 76.2 / 5.7 39.0 / 0.9 39.8 / 1.4 41.3 / 8.0 58.3 / 18.6
Llama-3.1-8B-Instruct 86.5 / 2.2 48.1 / 3.0 81.9 / 2.4 62.2 / 0.9 57.9 / 1.7 63.4 / 6.1 64.4 / 11.8
Qwen2.5-7B-Instruct 84.5 / 1.3 61.9 / 2.1 86.8 / 2.8 71.3 / 0.5 62.1 / 0.8 61.5 / 3.7 71.3 / 10.2
@@
Vicuna-7B 11.9 / 6.9 4.0 / 6.8 54.0 / 9.6 13.4 / 1.5 23.6 / 2.2 28.6 / 10.7 45.3 / 32.4
Llama-2-7B-Chat 25.8 / 4.7 5.4 / 6.0 60.0 / 6.6 16.4 / 1.2 24.8 / 1.8 26.2 / 10.5 48.2 / 19.2
Qwen1.5-7B-Chat 58.7 / 3.7 17.9 / 4.5 76.2 / 7.9 39.0 / 1.2 39.8 / 1.9 41.3 / 10.8 58.3 / 24.3
Llama-3.1-8B-Instruct 86.5 / 2.9 48.1 / 4.0 81.9 / 4.0 62.2 / 1.3 57.9 / 2.3 63.4 / 8.0 64.4 /15.2
Qwen2.5-7B-Instruct 84.5 / 1.7 61.9 / 2.6 86.8 / 3.7 71.3 / 0.7 62.1 / 1.1 61.5 / 4.7 71.3 / 13.0
@@
Llama-2-7B-Chat 25.8 / 2.5 5.4 / 3.4 60.0 / 3.9 16.4 / 7.3 24.8 / 1.1 26.2 / 5.5 48.2 / 11.2
Qwen1.5-7B-Chat 58.7 / 2.4 17.9 / 2.6 76.2 / 4.4 39.0 / 7.3 39.8 / 1.0 41.3 / 5.6 58.3 / 13.1
Llama-3.1-8B-Instruct 86.5 / 1.7 48.1 / 2.4 81.9 / 2.5 62.2 / 7.5 57.9 / 1.2 63.4 / 4.1 64.4 / 9.1
Qwen2.5-7B-Instruct 84.5 / 2.0 61.9 / 3.5 86.8 / 4.8 71.3 / 9.6 62.1 / 1.6 61.5 / 6.0 71.3 / 19.6
@@
Vicuna-7B 11.9 / 2.4 4.0 / 4.7 54.0 / 3.4 13.4 / 0.6 23.6 / 1.0 28.6 / 3.4 45.3 / 14.1
Llama-2-7B-Chat 25.8 / 2.0 5.4 / 3.9 60.0 / 2.9 16.4 / 0.4 24.8 / 0.7 26.2 / 3.8 48.2 / 11.4
Qwen1.5-7B-Chat 58.7 / 1.4 17.9 / 3.0 76.2 / 2.4 39.0 / 0.3 39.8 / 0.7 41.3 / 3.0 58.3 / 9.4
Llama-3.1-8B-Instruct 86.5 / 0.5 48.1 / 1.5 81.9 / 0.8 62.2 / 0.2 57.9 / 0.5 63.4 / 1.4 64.4 / 4.4
Qwen2.5-7B-Instruct 84.5 / 1.7 61.9 / 2.5 86.8 / 2.2 71.3 / 0.2 62.1 / 0.5 61.5 / 2.3 71.3 / 16.4
Table 14: Performance acc(%) / Activated proportion MUI(%) of the selected model. With activation-based importance (Equation 14) and top score ($k=0.95$) as the threshold.

### v3

@@ B.1 Utility Law: the Relationship Between MUI and Performance
Llama-2-7B-Chat 25.8 / 4.3 5.4 / 5.5 60.0 / 6.3 16.4 / 1.0 24.8 / 1.7 26.2 / 8.7 48.2 / 17.6
CodeLlama-7B-Instruct 23.0 / 5.1 6.0 / 8.0 49.5 / 9.6 34.1 / 2.0 44.7 / 3.3 21.8 / 8.5 41.4 / 26.3
Qwen2.5-7B-Instruct 84.5 / 2.3 61.9 / 3.5 86.8 / 5.4 71.3 / 0.9 62.1 / 1.6 66.0 / 6.6 71.3 / 17.7
Qwen2.5-Coder-7B-Instruct 79.6 / 4.3 49.5 / 7.5 83.8 / 11.1 75.0 / 1.8 65.9 / 3.0 57.8 / 13.1 65.6 / 34.4
Qwen2.5-Math-7B-Instruct 92.1 / 2.3 76.7 / 4.0 66.3 / 3.3 53.0 / 1.0 47.9 / 1.5 29.1 / 3.7 51.2 / 17.3
@@
Llama-2-7B-Chat 25.8 / 4.3 5.4 / 5.5 60.0 / 6.3 16.4 / 1.0 24.8 / 1.7 26.2 / 8.7 48.2 / 17.6
Llama-Code-Leakage 22.2 / 4.3 5.2 / 4.8 60.7 / 5.2 35.4 / 2.2 49.7 / 1.5 23.6 / 7.7 48.1 / 15.6
Llama-Math-Leakage 39.0 / 4.6 8.2 / 6.4 49.5 / 8.2 5.5 / 1.0 8.6 / 2.0 25.2 / 6.6 43.9 / 25.8
Llama-3.1-8B-Instruct 86.5 / 3.5 48.1 / 5.0 81.9 / 5.0 62.2 / 1.5 57.9 / 2.7 65.4 / 9.0 64.4 / 18.0
Llama3.1-Code-Leakage 86.6 / 4.0 40.2 / 5.5 81.2 / 6.2 81.7 / 1.6 69.1 / 2.4 48.1 / 8.2 66.3 / 18.5
Llama3.1-Math-Leakage 89.8 / 3.8 47.4 / 5.2 81.6 / 5.9 54.9 / 1.8 50.5 / 2.5 48.9 / 7.6 66.1 / 18.0
Qwen2.5-7B-Instruct 84.5 / 2.4 61.9 / 3.5 86.8 / 5.4 71.3 / 0.9 62.1 / 1.6 66.0 / 6.6 71.3 / 17.7
Qwen2.5-Code-Leakage 78.6 / 2.0 42.5 / 3.7 88.9 / 5.6 85.4 / 1.9 73.5 / 2.0 56.8 / 6.7 56.8 / 14.5
Qwen2.5-Math-Leakage 98.1 / 2.8 84.6 / 4.1 57.4 / 4.5 2.2 / 1.0 1.8 / 1.7 31.4 / 6.1 44.7 / 14.2
@@ B.6.3 Impact of Different Neuron Interpretation Methods
To demonstrate that our analysis is both generalizable and inclusive, we expand to incorporate various interpretation methods. Continuing from our previous threshold selections B.6.3, we conduct analyses using these methods to evaluate changes in the Model Utilization Index (MUI), with a particular focus on the consistency of MUI trends for different model. For some threshold values that do not integrate effectively, we randomly selected a threshold value within the testing range to facilitate a quantitative comparison. The results, which encompass the full parameter range, are detailed in the Figure 27. After applying different interpretation methods, we found that the MUI derived from various approaches (Table 13 and Table 14) correlates well with our experiment selections (Table 11), achieving an average Spearman correlation coefficient of 0.921. This indicates that the overall trend is relatively consistent that model capability and MUI is in a reverse Relationship. However, due to differences in the definition of importance, some models, such as Qwen2.5, exhibit a higher MUI when using activation-based methods compared to Llama3.1.
@@
Vicuna-7B 11.9 / 6.0 4.0 / 6.4 54.0 / 8.5 13.4 / 1.3 23.6 / 2.0 28.6 / 8.9 45.3 / 27.4
Llama-2-7B-Chat 25.8 / 4.3 5.4 / 5.5 60.0 / 6.3 16.4 / 1.0 24.8 / 1.7 26.2 / 8.7 48.2 / 17.6
Qwen1.5-7B-Chat 58.7 / 3.7 17.9 / 4.6 76.2 / 7.3 39.0 / 1.3 39.8 / 1.9 41.3 / 9.4 58.3 / 21.3
Llama-3.1-8B-Instruct 86.5 / 3.5 48.1 / 5.0 81.9 / 5.0 62.2 / 1.5 57.9 / 2.7 65.4 / 9.0 64.4 / 18.0
Gemma2-9B-Instruct 82.3 / 2.6 33.5 / 4.3 89.5 / 4.7 63.4 / 0.9 61.2 / 1.5 57.5 / 6.1 73.3 / 14.6
Qwen2.5-7B-Instruct 84.5 / 2.3 61.9 / 3.5 86.8 / 5.4 71.3 / 0.9 62.1 / 1.6 66.0 / 6.6 71.3 / 17.7
@@
Vicuna-7B 11.9 / 7.5 4.0 / 7.9 54.0 / 10.6 13.4 / 1.6 23.6 / 2.5 28.6 / 11.2 45.3 / 33.6
Llama-2-7B-Chat 25.8 / 5.3 5.4 / 6.8 60.0 / 7.7 16.4 / 1.3 24.8 / 2.1 26.2 / 10.9 48.2 / 21.3
Qwen1.5-7B-Chat 58.7 / 4.5 17.9 / 5.6 76.2 / 9.2 39.0 / 1.5 39.8 / 2.4 41.3 / 11.5 58.3 / 25.5
Llama-3.1-8B-Instruct 86.5 / 3.6 48.1 / 5.1 81.9 / 5.3 62.2 / 1.6 57.9 / 2.8 63.4 / 9.4 64.4 / 18.7
Qwen2.5-7B-Instruct 84.5 / 2.0 61.9 / 3.1 86.8 / 4.8 71.3 / 0.8 62.1 / 1.4 61.5 / 6.0 71.3 / 16.0
@@
Llama-2-7B-Chat 25.8 / 2.5 5.4 / 3.4 60.0 / 3.9 16.4 / 0.7 24.8 / 1.1 26.2 / 5.5 48.2 / 11.2
Qwen1.5-7B-Chat 58.7 / 2.4 17.9 / 2.6 76.2 / 4.4 39.0 / 0.7 39.8 / 1.0 41.3 / 5.6 58.3 / 13.1
Llama-3.1-8B-Instruct 86.5 / 1.7 48.1 / 2.4 81.9 / 2.5 62.2 / 0.8 57.9 / 1.2 63.4 / 4.1 64.4 / 9.1
Qwen2.5-7B-Instruct 84.5 / 2.0 61.9 / 3.5 86.8 / 4.8 71.3 / 0.9 62.1 / 1.6 61.5 / 6.0 71.3 / 19.6
@@
Vicuna-7B 11.9 / 2.4 4.0 / 3.7 54.0 / 3.4 13.4 / 0.6 23.6 / 1.0 28.6 / 3.5 45.3 / 13.6
Llama-2-7B-Chat 25.8 / 2.4 5.4 / 4.4 60.0 / 3.6 16.4 / 0.6 24.8 / 0.8 26.2 / 3.8 48.2 / 15.4
Qwen1.5-7B-Chat 58.7 / 1.9 17.9 / 3.5 76.2 / 2.6 39.0 / 0.4 39.8 / 0.8 41.3 / 3.7 58.3 / 10.1
Llama-3.1-8B-Instruct 86.5 / 0.9 48.1 / 2.0 81.9 / 1.2 62.2 / 0.3 57.9 / 0.6 63.4 / 2.9 64.4 / 6.5
Qwen2.5-7B-Instruct 84.5 / 1.8 61.9 / 4.0 86.8 / 2.8 71.3 / 0.4 62.1 / 0.8 61.5 / 3.8 71.3 / 14.0
Table 14: Performance acc(%) / Activated proportion MUI(%) of the selected model. With gradient-based importance (Equation 14) and top score ($k=0.95$) as the threshold.